OpenStudy (jiteshmeghwal9):

Solve the following equation by method of factor :- \[x^4+2x=3x^2\]

OpenStudy (anonymous):

First of all \[ x^4-3x^2+2x=0 \]

OpenStudy (anonymous):

Second of all \[ x(x^3-3x+2)=0 \]This means \(x=0\) is a solution. Now we just need to find solutions for: \[ x^3-3x+2=0 \]

OpenStudy (jiteshmeghwal9):

thanx i gt it :)

OpenStudy (anonymous):

You do?

OpenStudy (jiteshmeghwal9):

yup :)

OpenStudy (anonymous):

What are the remaining solutions?

OpenStudy (jiteshmeghwal9):

wait its 3 or 2

OpenStudy (jiteshmeghwal9):

x^3 or x^2

OpenStudy (jiteshmeghwal9):

@wio

OpenStudy (anonymous):

It is \(x^3\) not \(x^2\)

OpenStudy (anonymous):

Since \(x^4=x(x^3)\)

OpenStudy (jiteshmeghwal9):

oh u plz go on, i thought it is x^2

OpenStudy (anonymous):

Cubic equations have no simple solution. However there are guesses which are likely to work.

OpenStudy (anonymous):

The leading term is \(1x^3\) and the final term is \(2\). We take the coefficient of the leading and final term and find their factors. For the leading term it is just \(1\) For the final term, it is \(1,2\). If there is a rational root, it will be of the form \[ \pm \frac{1}{1,2} \]That is to say, the rational root, if there is one, would be \[ 1,-1,1/2,-1/2 \]

OpenStudy (anonymous):

So we can try these guesses here.

OpenStudy (anonymous):

Putting in \(1\) gives us \((1)^3-3(1)+2=1-3+2=0\) So fortunately, \(x=1\) is a root.

OpenStudy (anonymous):

This means we can factor \(x-1\).

OpenStudy (jiteshmeghwal9):

How can u say that the rational roots will be of the form \[\pm \frac{ 1 }{ 1,2 }?\]

OpenStudy (anonymous):

It's a particular theorem

OpenStudy (jiteshmeghwal9):

can u elaborate it ?

OpenStudy (anonymous):

Thing is, it might not even have rational roots.

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Rational_root_theorem

OpenStudy (jiteshmeghwal9):

Isn't there any other easier solution ?

OpenStudy (anonymous):

For cubic functions? You can guess and check. There are cubic formulas but they are extremely coplicated.

OpenStudy (anonymous):

Calculators will use newton's method to find solutions.

OpenStudy (anonymous):

Or bisection method.

OpenStudy (jiteshmeghwal9):

Ok thx for ur help & contribution :)